Meeting Information

  • 1600 UT (1000 MDT)
  • AOC 317; IP: 146.88.16.8; phone: +1 575-835-7370

Agenda/Minutes

News, Travel, Visitors (McMullin)

  • Overall status: Much RSRO observing using current observing recommendations and older data acquisition software; some issues with missing/mislabeled scans (cover in RSRO discussion). Expecting new version of CBE/CM this week as well as correction to BlB configuration problems.
  • Key dates:
    • 13 Sep: Begin move to DnC configuration
    • 15 Sep: Call for Proposals
      • Note: Ku/X-wide will not be offered in main call (allowed for exploratory).
    • 01 Oct: Proposal deadline (A config; lowband for end of B config)
    • 04 Oct: Begin move to C configuration
  • Staff Travel:
    • Hayward: 10 Aug - 30 Aug
    • Chandler: 01 Sep - 15 Sep
    • DRAO: Carlson: 16 Jul - 19 Aug
    • Rupen: 11 Sep - 21 Sep
    • Sowinski: 20 Sep - 15 Oct
    • McMullin: 01-06 Oct
    • van Moorsel: 11-15 Oct
    • Wrobel: 24 Oct-27 Oct
    • DRAO: Dougherty: 11 Oct - 12 Oct (SKA; ABQ)
  • Visitors:
    • Bob Sault: 26 July - 25 August
    • Brent Carlson: 18 October - 25 October; second trip post SKA meeting TBD
    • DRAO: Dougherty: 12 Oct - 17 Oct (NRAO-DSOC)
    • Adam Leroy/Joe Lazio: 13-15 Oct (phased array testing)

Hardware status and updates (Langley, Perley, Rupen)

  • Array (Perley/Langley):
    • Notes: No big news; receiver flow continues; EA08 issues not understood but it's back in the array currently. EA02 IF C still has 125 kHz spikes (under investigation); EA11 RR has issues (Perez); EA27 RR amp jumps (Perez). Note from Ken that there was a power glitch on Monday; 3 antennas had issues and weren't immediately recovered.
    • Links:
  • WIDAR (Rupen):
    • Notes: Final(?) BlB delivery made; hardware at the site. Final switch to handle all the BlB has been ordered.

Software status and updates (Butler)

  • Notes: Anticipating new version of CBE software that resolves most of missing scan issues; outstanding issues with Executor tardiness leading to mislabeled scans; additional logging added but no catastrophic failures since the change to sift through. Archive access speeds affected by VO services which have been temporarily turned off. CASA is going through a management change beginning Oct (Jeff Kern will take over for Nick) at the 10% level and ramping up to 100% by January due to existing ALMA obligations. JO: CASA development progress meeting this week for the October release (things are looking reasonable). OPT: Bug fix update at 1000 MDT today-delayed to Thursday; (brief outage; release notes at: https://staff.nrao.edu/evla/sss/opt/sssOptReleases.shtml). Executor crash last night seems to be related to the version of JAXB; looking to update (Clark); new version of CBE under trial but shows great improvement in dropped/missing scans. Next focus will be on averaging. MR: Currently a problem with the BlB configuration (dropping frames to the CBE - all data is lost); being worked on. Discussion on how to send out information to users on some of the detailed problems; done case-by-case. JO: CASA has an averaging bug that is being investigated but is a show-stopper in some cases.
  • Links:

Observing recommendations update (Ops)

  • Notes: Pending deployment of new software; not tested over the weekend so full measure not available until next week.
  • Link: https://staff.nrao.edu/wiki/bin/view/EVLA/RSROObservingPreparationGuidelines
  • Proposal (to handle interim issues)::
    • Recommend putting critical calibration in SBs twice (beginning&end, beginning& middle, etc).
    • Recommend submitting a short (30 min) SB that contains critical calibration (bandpass/flux scale) that can be run independently from the science SB.

Commissioning priorities for the week (McMullin, Chandler)

  • 1) RSRO/ECSO science
  • 2) 3-bit sampler testing (Wednesday & Friday; possible Thursday)
  • 3) OSRO/RSRO blockers (focused on missing data/data quality issues)
  • 4) CSV testing
  • 5) Demo science observing
  • 6) OSRO science observing

Commissioning Reports.

Polarization: Taylor: C band results (with RFI sweep).

Spectral Line: Ott: Doppler setting recommendations/status.

  • CASA defect has prevented key analysis; more to come.

RFI measurements: Momjian: High/low frequency band sweeps. Link: http://science.nrao.edu/evla/observing/RFI/

Switched power update (Butler/Rupen)

  • Myers; can be plotted with script. concentrator had caused problems but has been resolved; will run more this week. Issue of handling in the ASDM is still being worked on - Cal_Device table use, etc.

In preparation

Polarization: Multi-band leakage/PA (Myers, Marvil): Needed.

Spectral Line: Zeeman calibration (Momjian); H2O/CH3OH masers on known sources (simple V); Observations completed.

Commissioning Tests

Calibration: Dhawan: Frequency offset tests and gain stability

Polarization: Antenna cross polarization (Perley; Need script.); LSCX on 3C147; Tuesday-Wednesday (0800-0900 LST); Delayed

Spectral Line: Bandpass stability,multi-band on 3C84 et al. (Goss/Chomiuk); TBS

Other:

RSRO/ECSO: Reports (Chandler, Kepley, Heesen, Green; Momjian, Owen, Wrobel/Ott, Marvil)

Project Band(s) Test Code Test Status Allocation Status (D) Notes JIRA
RSRO            
AA30 Q TRSR0003 - 0/51 Ready
1773293 (5.5 hr; 0630-0830)
RSRO-1
AB1345 L TRSR0004 21May10 23/40 In Progress;Needs SB
1659581 (7 hr;1430-1530)
1747558 (8 hr;1430-1530)
1800808 (8 hr; 1430-1530)
RSRO-2
AB1346 K TRSR0005
TSPE0001
23Jun10 0/42 Ready
TSPE0001 1177894 (2 hr; 1700-1800)
1805547 (3 hr; 1630-1800)
RSRO-3
AC982 CKKaQ TRSR0002 28Jul10 22/30 In Progress
1537991 (9 hr; 0000-0000)
1567773 (2 hr;0230-0230)
1553081 (2 hr;0030-0030)
1761189 (2 hr; 0030-0030)
1769750 (2 hr; 0030-0030)
1545148 (4 hr; 2330-2330)
1558397 (4.5 hr; 2330-0001)
RSRO-4
AH1006 LCXKa TRSR0006 - 16/36 In Progress; Need SBs
1787521 (4 hr;2100-0200)
1800294 (4 hr; 2100-0200)
1807363 (8 hr; 2100-0200)
RSRO-5
AK726 Ka TRSR0007 - 2/43 In Progress; Need SBs
1763407 (2 hr; 2200-0100)
RSRO-6
AL746 CKa TRSR0008 05Jul10 6.5/11 In Progress
1406401 (1.5 hr; 2100-2330)
1783341 (2 hr; 1500-1830)
1775937 (1.5 hr; 2030-2330)
1777156 (1.5 hr;0130-0400)
1779817 (2.5 hr;1000-1230)
RSRO-7
AM1014 C TRSR0009 22Jun10 0/0 C config RSRO-9
AS1015 C TRSR0010 21May10 4/4 Complete
1482295 (1 hr; 0830-1630),1626830 (1 hr; 0930-1030)
RSRO-11
AT374 C TRSR0011 25Jul10 0/0 C config RSRO-12
AV318 L TRSR0012 - 12/12 Complete; PI converted to OSRO
1491201 (6 hr; 0930-1030)
RSRO-13
ECSO          
AW775 L TRSR0017 05Jul10 40/72 In Progress
1660006 (8)
-
AM1005 KKaQ TRSR0014 - 3/20 In Progress; Needs SB
1748530 (3 hr; 1830-1830)
-
AM1038 Q TRSR0019 21May10 0/10 Ready
1491016 (5 hr; 0430-0730)
1638323 (5 hr; 0730-0930)
1731355 (5 hr; 0030-0200)
1735390 (5 hr;0200-0330)
-
AM1040 Ka TRSR0019 21May10 10/10 Complete
1736712 (5 hr; 0430-0730)
1737567 (5 hr; 0730-0930)
1738422-linked (5 hr; 0030-0200)
1739296 (5 hr;0200-0330)
-
AC992 L TRSR0018 26Jun10 4/4 Complete
1426692
1427000
1427915
1705655
-
AO255 Ka TRSR0016 - 16/40 Ready
1423157 (4 hr;1400-1430)
1466135 (4 hr; 1230-1330)
1467049 (4 hr;1500-1730)
1467331 (8 hr;1230-1230)
1481381 (8hr;1400-1430)
1756969 (4 hr;1500-1730)
-
AM1018 Ka TRSR0015 - 0/48 No SBs submitted -
AB1353 Ka TRSR0013 - 0/51 No SBs submitted -

  • Kepley: data looks fine; did lose a reference pointing; more SBs by Friday. AH1006 - weak detection in Stokes Q; problems with split/flagging - Myers couldn't reproduce.AO255; Frazer can see the line; needs more integration time. AM1018 - want to design a test to evaluate OTF mapping.
    • Resident Schedule: Posted on home page of twiki now!
      • 01 Jan - 16 May: Adam Deller (AC982: RSRO-4): 300A
      • 31 Jan - 31 Jul: Laura Chomiuk (AS1015/AK726: RSRO-11/RSRO-6): 200C
      • 04 Feb - 01 Jun: Jacqueline van Gorkom (AV318: RSRO-13): 200D
      • 25 Feb - 01 Jun: James Miller-Jones (AM1014: RSRO-9): 200E
      • 02 Mar - 07 Jun: Manuel Aravena (AA330: RSRO-1): 300G
      • 16 Mar - 08 Apr: Dave Green (AB1345: RSRO-2): 300G
      • 20 Apr - 30 Jun: Crystal Brogan (AB1346: RSRO-3): 200B
      • 20 Apr - 30 Jun: Todd Hunter (AB1346: RSRO-3): 200B
      • 30 Apr - 15 Jul: Joseph Lazio (AC982: RSRO-4): 266
      • 07 Jun - 16 Sep: Laura Perez (AC982: RSRO-4): 200E
      • 28 Jun - 05 Oct: Volker Heesen (AH1006: RSRO-5): 200G
      • 05 Jul - 26 Oct: Russ Taylor (AT374: RSRO-12): 301
      • 21 Jul - 23 Sep: Dave Green (AB1345: RSRO-2): 300G
      • 01 Aug - 31 Aug: Peter Hofner (AH1011/10B-124): 200G/7496
      • 01 Sep - 30 Nov: Satoko Takahashi (AT380/10B-126): 200D/7098
      • 01 Sep - 30 Sep: Ken Kellerman (AK736/10B-104): 322/7373
      • 03 Sep - 22 Dec: Gregg Hallinan (AH1023/10B-224): 200B/7334
      • 15 Sep - 15 Dec: Adam Leroy (AL746: RSRO-7): 200G/7496
      • 01 Oct - 01 Jan: Hendrik Linz (AC982: RSRO-4): 200C
      • 01 Jan - 01 Apr: Jane Greaves (AC982: RSRO-4): 300G
      • 02 Jan - 02 Apr: Stuart Corder (AC982: RSRO-4): 300G
    • Current Residents
      • Laura Perez (POC: Claire Chandler); area(s): High frequency observing
      • Dave Green (POC: Sanjay Bhatnagar): area(s): algorithm tests for wideband imaging (currently away).
      • Volker Heesen (POC: Michael Rupen); area(s): polarimetry
      • Peter Hofner (POC: TBD); area(s): post-processing
      • Russ Taylor (POC: Rick Perley); area(s): polarimetry

Demo Science (Chandler/Condon): Project Codes: *TDEM*

    • Status: Begun; list requires prioritization; limited remaining time in D-config; Will try for IRC+10216 this week as possible.
    • Testing
Project Status Time Comments SB
Crab nebula Complete 1x10 L-band OSRO1/OSRO2 TDEM0007_sb1524351 (10 hr: 0030-1030 LST)
IRC+10216 Ready (HF) 33 K, Ka, Q spectral sweep; no recirculation; pending test TDEM0003_sb1654086 (30 min: 0730-1130 LST); RSRO 1 GHz, dual pol; 125 kHz
1525160 (3 hr;0730-1130); OSRO 8 MHz BW (HC7N/HC5N)
1542567 (3 hr; 0730-1130); OSRO 8 MHz BW (HC5N/C3N)
1542814 (3 hr; 0730-1130 LST); OSRO 8 MHz BW (C3N)
1543061 (3 hr; 0730-1130); OSRO 8 MHz BW (C4H)
M87 Ready (HF) 1x6 Ka TDEM0006_sb1470572 (not submitted; 6 hr: 0900-1000 LST)
M51 Ready (MR) 1x8 L band; needs corr test TDEM0008_sb1542082 (7.75 hr: 0830-1130 LST)
Cygnus A Ready 3x6 1x10 possible; 12 OSRO hours obtained Withdrawn
M82 Complete 1x2 L band TDEM0010_sb1660855 (2 hr: 0400-0930 LST)
Herc A Ready 1x5 CX bands TDEM0011_sb1729424 (5 hr; 1500-1700)

Commissioning Group updates

AOB

Misc Actions

  • [McMullin/Frail]: %REDWrite up disposition of RR items.
  • [McMullin/Butler/Rupen/Owen]: Resolve discussion of requirements for DnC test of 4-band.
  • [McMullin]: Confirm restriction of RSRO scan length at 90 s
  • [Owen]: Outline issue with 90s scan length restriction.
  • [McMullin/Durand]: Clean up outdated MainSaver issues.
  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley/McMullin]: SPW/IF/BB lexicon and mapper
  • List of known problems (can be used by OSRO commissioning team (van Moorsel)
      • Med Priority [Rupen]: X, Ku band subbands are inverted (IF 1 is labelled as IF4, IF2 as 3, IF3 as 2, IF 4 as 1). Noted that this had been tested but there were enough changes to warrant a re-do.
      • Med Priority [TBD]: Phase offsets within scans (phase jump at scan boundaries). R&L phase offset on some scans - probably L302 LO problem, needs tracking down. MR: Seen on EA07, EA25 (both RR and LL)
      • Med Priority [TBD]: Added. Phase change when frequency changes (needs to be tested).
      • Med Priority [TBD]: Anti-aliasing does not behave as advertised.
      • Low Priority [TBD]: First scan often bad.
      • Low Priority [TBD]: Phase (and amp) wobbles; long history.
-- JosephMcMullin - 23 Aug 2010
Topic revision: r6 - 2010-08-29, jmcmulli
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback